# Open-Set Domain Adaptation Under Background Distribution Shift: Challenges and A Provably Efficient Solution

### v2

@@ 5.2 A simple Extension of Constrained Learning for OSDA
To account for the known classes $Y=1,\ldots,k$ and $L$ novelty heads, we construct $w=[w^{c},w^{\alpha}]$, such that $w:{\mathbb{R}}^{d}\rightarrow{\mathbb{R}}^{L+k}$. $w^{\alpha}$ is responsible for detecting novelties with various proportions while $w^{c}$ determines the known class of the input $\mathbf{x}$}. See algorithm˜1 for further details. Note that training multiple novelty heads ($w^{\alpha}_{1},w^{\alpha}_{2},...,w^{\alpha}_{L}$]) simultaneously is much more computationally efficient than training a separate model for each candidate $\hat{\alpha}$ in a loop. We hypothesize that having a shared representation for both tasks enables the model to learn a simpler representation that fits both tasks, yielding favorable generalization bounds as suggested by theory on multitask learning (baxter_model_2000; maurer_benet_16).
@@
In our experiments, the representation $\phi({\mathbf{x}})$ is either trained from scratch or built on a pretrained backbone with added fully connected layers. For small datasets (e.g., CIFAR100), we train $\phi(x)$ as ResNet18 (he_deep_2016) from scratch. For larger ones (e.g., SUN397), we use pretrained $\phi(x)$ such as ResNet50 (ImageNet1K_V1) (Russakovsky2015imagenet) or a CLIP-pretrained ViT-L/14 encoder (radford2021clip). For text datasets (e.g., Amazon Reviews), we use RoBERTa (liu_roberta_2019) for embeddings. See section˜A.5 for details.
@@ 6.2 Datasets
Most of the large scale image classification models are trained on ImageNet dataset. Hence we use a similar large scale dataset, SUN397 (jianxiong2010sun397) having completely different categories than ImageNet. It is a popular scene understanding dataset with a three-level hierarchy of scene categories jianxiong2010sun397, exploiting this structure to create distribution shifts by varying subtype ($y$) proportions within primary class labels $\mathcal{Y}$. Indoor scenes (e.g., shopping/dining places, workplaces) serve as known (in-distribution) classes, while outdoor scenes are randomly selected as novel classes. Varying subtype proportions in indoor categories causes a background shift since $\mathrm{Supp}(P_{{\mathcal{T}},[k]})\subseteq\mathrm{Supp}(P_{{\mathcal{S}}})$. We also include a Amazon Product Reviews text dataset (ni_justifying_2019) to demonstrate the versatility of the method across multiple modalities. Classes are different product categories (prime pantry, musical instruments, etc.), and induce background/covariate shift in known classes based on positive vs. negative sentiment in the review. The novel class is an unknown product category. Finally, we include small scale CIFAR100 (Krizhevsky09learningmultiplecifar) dataset to evaluate the adaptive methods (like DD, nnPU, uPU, BODA, OSDA) particularly in scenarios when their feature extractors are trained from scratch. Similar to the SUN397 dataset, we leverage the inherent hierarchies of CIFAR100 classes to create a natural background shift based on subtypes (e.g. dolphins) of known primary categories like aquatic mammals. Novel classes are chosen from other categories in distinct branches of the class hierarchy. Table˜1 provides a summary of characteristics and further details on each dataset are in section˜A.5.1.
@@
Evaluation metrics: We primarily use Area Under ROC Curve (AUROC) and Area Under Precision-Recall Curve (AUPRC) to evaluate the novel category detection performance, while we use Open-Set Classification Rate (OSCR) to summarize overall OSDA performance for all the methods  [^8] Section A.4.1 in appendix provides detailed argument for the preference of AUPRC over AUROC mainly when minority class proportions are very low.. OSCR measures the trade-off between correct classification rate of the known classes and false positive rate of the novel samples, we refer to dhamija2018oscr for further details about the metric.
@@ 6.3 Baseline methods
We include adaptive methods from novelty detection that access both labelled source and unlabelled target data. These baselines include domain discriminator DD, (elkan2008learning; duplessis2014analysis), uPU (duplessis2014analysis) and nnPU (kiryo2017positive). These methods are modified for OSDA through joint learning approach enabled by a simple architectural modification like figure 2 and aggregating the loss components from both the tasks. We also include another popular OSDA baselines BODA (saito2018open) and PULSE (garg22adaptation) that are agnostic to input data modality. Another adaptive baseline is wuyang2023anna, however, this is specifically designed for vision data. Hence, we report its performance on SUN397 dataset in table˜3 along with other vision-specific baselines.
We include an entropy-based method, ARPL (Chen_2021) with Maximum Logit Score as proposed in (vaze2022openset). Such methods are non-adaptive as they do not access target data, yet are a popular choice for novelty detection. [^9] We are using ARPL on SUN397. The paper that proposed ARPL also introduced an improvement with Confused Sampling (ARPL+CS) that uses GANs, but that proved challenging to apply in the SUN397 dataset We also tested simple and popular baselines such as MSP hendrycks2016baseline and results can be found in appendix tables 15, 16. We further include SHOT (liang2021reallyneedaccesssource), CAC (Miller2021ClassAC) and a zero-shot OOD detection method (ZOC) proposed in Esmaeilpour_Liu_Robertson_Shu_2022. Further details about training and hyperparameters are in Appendix A.5.2.
@@ 6.5 Discussion
From tables 3, 8 and 9 we observe that all the methods that involve training a closed-set classifier suffer from background shift. Hence, background shift not only harms closed-set performance but also that of open-set recognition. If we had a better closed set classifier, in the sense that it was more robust to the background shift, then we probably would’ve ended up with better OSDA performance too. More specifically, CoLOR seems to primarily help when standard techniques like long training, label smoothing, and other techniques used in vaze2022openset are insufficient to train a model that’s robust to shift. ZOC performs zero-shot open-set classification using pretrained CLIP and text decoder models without accessing source and target data. Hence, it is not impacted by background shift as expected. We have provided further results on CIFAR100 dataset in 10 & 11 although CIFAR100 is a much simpler dataset and hence we see near perfect scores by ZOC due to the use of large CLIP ViT-B/32 model pretrained on huge image caption datasets. We see in tables 8 & 9 that non-adaptive methods do not perform well when scaled to larger SUN397 dataset.
@@ 7 Future Work
Future research could focus on refining the search criteria for the optimal model head by selecting an appropriate $\beta$ threshold. Additionally, a robust theoretical underpinning is needed to explain why shared representations improve the classification of known classes under distribution shifts. Integrating test-time or model-free domain adaptation methods, such as those by wang_tent_2020, zhang_model_free_2023, or saito2018open, could further enhance OSDA performance. It would be valuable to compare our approach with large foundation models. Although curating such a benchmark would require access to publicly available pretraining datasets like LAION, which would require significant effort and present unforeseen challenges.
@@ Acknowledgments
Shravan S. Chaudhari and Rama Chellappa acknowledge support from the DARPA TIAMAT program under Grant No. HR00112490422. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the U.S. Government.
@@ A.1 Theoretical analysis
For a Gaussian two-feature PU problem, let $\hat{\eta}=\eta/\|\eta\|$ and $\hat{\mu}$ accordingly. We have that the AU-ROC of a model ${\mathbf{w}}\in{\mathbb{R}^{d}}$ is $Q(\frac{\langle{\mathbf{w}},\eta-\mu\rangle}{\sigma\sqrt{(2\|{\mathbf{w}}\|^{2}-\langle{\mathbf{w}},\hat{\eta}\rangle^{2}-\langle{\mathbf{w}},\hat{\mu}\rangle^{2})}})$.
@@
$=P(\langle{\mathbf{w}},-\eta+\xi_{\mathrm{novel}}\rangle<\langle{\mathbf{w}},-\mu+\xi_{\mathrm{{\mathcal{T}},0}}\rangle)$
$=P(\langle{\mathbf{w}},-\eta+\mu\rangle<\langle{\mathbf{w}},-\xi_{\mathrm{novel}}+\xi_{\mathrm{{\mathcal{T}},0}}\rangle),$
where $\xi_{\mathrm{novel}}\sim\mathcal{N}(0,\sigma^{2}(I_{d}-\hat{\mu}\hat{\mu}^{\top}))$ and $\xi_{{\mathcal{T}},0}\sim\mathcal{N}(0,\sigma^{2}(I_{d}-\hat{\eta}\hat{\eta}^{\top}))$. Then we have that $\langle{\mathbf{w}},\xi_{\mathrm{novel}}-\xi_{{\mathcal{T}},0}\rangle\sim\mathcal{N}\left(0,\sigma^{2}\left(2\|{\mathbf{w}}\|^{2}-\langle{\mathbf{w}},\hat{\eta}\rangle^{2}-\langle{\mathbf{w}},\hat{\mu}\rangle^{2}\right)\right)$. Dividing both sides of the argument above by a factor $\sigma\sqrt{2\|w\|^{2}-\langle{\mathbf{w}},\hat{\eta}\rangle^{2}-\langle{\mathbf{w}},\hat{\mu}\rangle^{2}}$ we get that for a variable $\xi\sim\mathcal{N}(0,1)$
$\mathrm{AU-ROC}({\mathbf{w}})$ $=P\left(\frac{\langle{\mathbf{w}},\mu-\eta\rangle}{\sigma\sqrt{2\|{\mathbf{w}}\|^{2}-\langle{\mathbf{w}},\hat{\eta}\rangle^{2}-\langle w,\hat{\mu}\rangle^{2}}}<\xi\right)$
$=Q\left(\frac{\langle{\mathbf{w}},\mu-\eta\rangle}{\sigma\sqrt{2\|{\mathbf{w}}\|^{2}-\langle{\mathbf{w}},\hat{\eta}\rangle^{2}-\langle{\mathbf{w}},\hat{\mu}\rangle^{2}}}\right)$
Note that when $\langle{\mathbf{w}},\mu-\eta\rangle>0$ then the AU-ROC is lower than $0.5$, and when $\langle{\mathbf{w}},\mu-\eta\rangle<0$ then $Q\left(\frac{\langle{\mathbf{w}},\mu-\eta\rangle}{\sigma\sqrt{2\|{\mathbf{w}}\|^{2}-\langle{\mathbf{w}},\hat{\eta}\rangle^{2}-\langle{\mathbf{w}},\hat{\mu}\rangle^{2}}}\right)\geq Q\left(\frac{\langle{\mathbf{w}},\mu-\eta\rangle}{\sqrt{2}\sigma\|{\mathbf{w}}\|}\right)$. Hence the RHS is a lower bound on the AU-ROC.
@@
$C_{d}\log(\frac{c_{d}}{\delta})\leq\frac{\sqrt{d}}{4\sqrt{N_{{\mathcal{S}}}+N_{{\mathcal{T}}}}},\frac{1}{2\sqrt{N_{{\mathcal{S}}}+N_{{\mathcal{T}}}}}\leq r_{\mu},r_{\eta}\leq\frac{r_{\mu}}{4+8r_{\mu}\sqrt{N_{{\mathcal{T}},1}}}-\frac{1}{\sqrt{N_{{\mathcal{S}}}+N_{{\mathcal{T}}}}}$
There exists $C_{d},c_{d}$ such that for any $\delta$, if we fix $\mu,\eta,r_{\mu},r_{\eta},N_{{\mathcal{S}}},N_{{\mathcal{T}},0},N_{{\mathcal{T}},1}$ that are in the above set, and drawing training data $(S_{\mathcal{S}},S_{\mathcal{T}})$ as described in problem definition. Let ${\mathbf{w}}_{\mathrm{\text{DD}}}$ as defined in eq.˜2, then with probability at least $1-\delta$ we have that $AU-ROC({\mathbf{w}}_{\mathrm{\text{DD}}})\leq 0.5$.
@@
$\mathcal{L}({\mathbf{w}},\lambda,\nu)=\|{\mathbf{w}}\|^{2}+\lambda_{{\mathcal{S}}}^{\top}(\mathbf{1}_{{\mathcal{S}}}-\bm{X}_{{\mathcal{S}}}{\mathbf{w}})+\lambda_{{\mathcal{T}}}^{\top}(\mathbf{1}_{{\mathcal{T}}}+\bm{X}_{{\mathcal{T}}}{\mathbf{w}})+\nu\cdot({\mathbf{w}}^{\top}(\mu-\eta))$
@@
$\nabla_{{\mathbf{w}}}\mathcal{L}({\mathbf{w}},\lambda,\nu)=2{\mathbf{w}}-\bm{X}^{\top}_{{\mathcal{S}}}\lambda_{{\mathcal{S}}}+\bm{X}^{\top}_{{\mathcal{T}}}\lambda_{{\mathcal{T}}}+\nu\cdot(\mu-\eta)=0$
$\Leftrightarrow{\mathbf{w}}=\frac{1}{2}\left(\nu\cdot(\eta-\mu)+\bm{X}^{\top}_{{\mathcal{S}}}\lambda_{{\mathcal{S}}}-\bm{X}^{\top}_{{\mathcal{T}}}\lambda_{{\mathcal{T}}}\right)$
@@
$\mathcal{L}(\lambda,\nu)=\lambda_{{\mathcal{S}}}^{\top}\mathbf{1}_{{\mathcal{S}}}+\lambda_{{\mathcal{T}}}^{\top}\mathbf{1}_{{\mathcal{T}}}-\frac{1}{4}\|\nu\cdot(\eta-\mu)+\bm{Z}^{\top}\lambda\|^{2}$
@@
Let $\alpha>0$, to be determined later, $\nu=\frac{(|U|-k)2\alpha}{|U|}$, and define the set $U=\mathcal{I}_{{\mathcal{T}},1}\cup\mathcal{I}_{\mathrm{Rand}}$ where $\mathcal{I}_{\mathrm{Rand}}$ has $\frac{|U|}{2\alpha}\nu$ indices randomly drawn from $\mathcal{I}_{{\mathcal{S}}}$ and half from $\mathcal{I}_{{\mathcal{T}},0}$ (not we should also set $|U|$ later). Set $\lambda=\frac{\alpha}{|U|}\mathbf{1}_{U}$ and plug back into the Lagrangian. We have that
$\bm{Z}^{\top}\lambda=\nu\mu+\frac{\alpha}{|U|}\left[k\eta+\sum_{i\in{U}}{\xi_{i}y_{i}}\right].$
@@
$\mathcal{L}(\alpha,\nu)=\alpha-\frac{1}{4}\|\eta(\nu+k\frac{\alpha}{|U|})+\alpha\bar{\xi}_{U}\|^{2}$
@@
$\mathcal{L}(\alpha)=\alpha-\frac{1}{4}\|\eta(\frac{\alpha(2|U|-k)}{|U|})+\alpha\bar{\xi}_{U}\|^{2}$
@@
$\|{\mathbf{w}}^{*}\|^{2}\geq\mathcal{L}(\alpha^{*})$ $=4\|\eta(\frac{2|U|-k}{|U|})+\bar{\xi}_{U}\|^{-2}$
$=4\left(r^{2}_{\eta}(2-k/|U|)^{2}+\|\bar{\xi}_{U}\|^{2}+2(2-k/|U|)\eta^{\top}\bar{\xi}_{U}\right)^{-1}$
@@
${\mathbf{w}}_{\mathrm{\text{DD}}}=2\left(r^{2}_{\mu}-\sigma r_{\mu}t\right)^{-1}\cdot\mu-4\sigma^{-2}kd^{-1}\cdot\bar{\xi}_{{\mathcal{T}},1},$
satisfies $y_{i}{\mathbf{w}}_{\mathrm{\text{DD}}}^{\top}{\mathbf{x}}_{i}\geq 1$ for all $i\in{U}$ and ${\mathbf{w}}_{\mathrm{\text{DD}}}^{\top}(\mu-\eta)\geq 0$.

### v3

@@ 5.2 A simple Extension of Constrained Learning for OSDA
To account for the known classes $Y=1,\ldots,k$ and $L$ novelty heads, we construct $w=[w^{c},w^{\alpha}]$, such that $w:{\mathbb{R}}^{d}\rightarrow{\mathbb{R}}^{L+k}$. $w^{\alpha}$ is responsible for detecting novelties with various proportions while $w^{c}$ determines the known class of the input $\mathbf{x}$}. See algorithm˜1 for further details. Note that training multiple novelty heads ($w^{\alpha}_{1},w^{\alpha}_{2},...,w^{\alpha}_{L}$]) simultaneously is much more computationally efficient than training a separate model for each candidate $\hat{\alpha}$ in a loop. We hypothesize that having a shared representation for both tasks enables the model to learn a simpler representation that fits both tasks, yielding favorable generalization bounds as suggested by theory on multitask learning (Baxter, 2000; Maurer et al., 2016).
@@
In our experiments, the representation $\phi({\mathbf{x}})$ is either trained from scratch or built on a pretrained backbone with added fully connected layers. For small datasets (e.g., CIFAR100), we train $\phi(x)$ as ResNet18 (He et al., 2016) from scratch. For larger ones (e.g., SUN397), we use pretrained $\phi(x)$ such as ResNet50 (ImageNet1K_V1) (Russakovsky et al., 2015) or a CLIP-pretrained ViT-L/14 encoder (Radford et al., 2021). For text datasets (e.g., Amazon Reviews), we use RoBERTa (Liu et al., 2019) for embeddings. See section˜A.5 for details.
@@ 6.2 Datasets
Most of the large scale image classification models are trained on ImageNet dataset. Hence we use a similar large scale dataset, SUN397 (Xiao et al., 2010) having completely different categories than ImageNet. It is a popular scene understanding dataset with a three-level hierarchy of scene categories Xiao et al. (2010), exploiting this structure to create distribution shifts by varying subtype ($y$) proportions within primary class labels $\mathcal{Y}$. Indoor scenes (e.g., shopping/dining places, workplaces) serve as known (in-distribution) classes, while outdoor scenes are randomly selected as novel classes. Varying subtype proportions in indoor categories causes a background shift since $\mathrm{Supp}(P_{{\mathcal{T}},[k]})\subseteq\mathrm{Supp}(P_{{\mathcal{S}}})$. We also include a Amazon Product Reviews text dataset (Ni et al., 2019) to demonstrate the versatility of the method across multiple modalities. Classes are different product categories (prime pantry, musical instruments, etc.), and induce background/covariate shift in known classes based on positive vs. negative sentiment in the review. The novel class is an unknown product category. Finally, we include small scale CIFAR100 (Krizhevsky, 2009) dataset to evaluate the adaptive methods (like DD, nnPU, uPU, BODA, OSDA) particularly in scenarios when their feature extractors are trained from scratch. Similar to the SUN397 dataset, we leverage the inherent hierarchies of CIFAR100 classes to create a natural background shift based on subtypes (e.g. dolphins) of known primary categories like aquatic mammals. Novel classes are chosen from other categories in distinct branches of the class hierarchy. Table˜1 provides a summary of characteristics and further details on each dataset are in section˜A.5.1.
@@
Evaluation metrics: We primarily use Area Under ROC Curve (AUROC) and Area Under Precision-Recall Curve (AUPRC) to evaluate the novel category detection performance, while we use Open-Set Classification Rate (OSCR) to summarize overall OSDA performance for all the methods  [^8] Section A.4.1 in appendix provides detailed argument for the preference of AUPRC over AUROC mainly when minority class proportions are very low.. OSCR measures the trade-off between correct classification rate of the known classes and false positive rate of the novel samples, we refer to Dhamija et al. (2018) for further details about the metric.
@@ 6.3 Baseline methods
We include adaptive methods from novelty detection that access both labelled source and unlabelled target data. These baselines include domain discriminator DD, (Elkan and Noto, 2008; du Plessis et al., 2014), uPU (du Plessis et al., 2014) and nnPU (Kiryo et al., 2017). These methods are modified for OSDA through joint learning approach enabled by a simple architectural modification like figure 2 and aggregating the loss components from both the tasks. We also include another popular OSDA baselines BODA (Saito et al., 2018) and PULSE (Garg et al., 2022) that are agnostic to input data modality. Another adaptive baseline is Li et al. (2023), however, this is specifically designed for vision data. Hence, we report its performance on SUN397 dataset in table˜3 along with other vision-specific baselines.
We include an entropy-based method, ARPL (Chen et al., 2021) with Maximum Logit Score as proposed in (Vaze et al., 2022). Such methods are non-adaptive as they do not access target data, yet are a popular choice for novelty detection. [^9] We are using ARPL on SUN397. The paper that proposed ARPL also introduced an improvement with Confused Sampling (ARPL+CS) that uses GANs, but that proved challenging to apply in the SUN397 dataset We also tested simple and popular baselines such as MSP Hendrycks and Gimpel (2016) and results can be found in appendix tables 15, 16. We further include SHOT (Liang et al., 2021), CAC (Miller et al., 2021) and a zero-shot OOD detection method (ZOC) proposed in Esmaeilpour et al. (2022). Further details about training and hyperparameters are in Appendix A.5.2.
@@ 6.5 Discussion
From tables 3, 8 and 9 we observe that all the methods that involve training a closed-set classifier suffer from background shift. Hence, background shift not only harms closed-set performance but also that of open-set recognition. If we had a better closed set classifier, in the sense that it was more robust to the background shift, then we probably would’ve ended up with better OSDA performance too. More specifically, CoLOR seems to primarily help when standard techniques like long training, label smoothing, and other techniques used in Vaze et al. (2022) are insufficient to train a model that’s robust to shift. ZOC performs zero-shot open-set classification using pretrained CLIP and text decoder models without accessing source and target data. Hence, it is not impacted by background shift as expected. We have provided further results on CIFAR100 dataset in 10 & 11 although CIFAR100 is a much simpler dataset and hence we see near perfect scores by ZOC due to the use of large CLIP ViT-B/32 model pretrained on huge image caption datasets. We see in tables 8 & 9 that non-adaptive methods do not perform well when scaled to larger SUN397 dataset.
@@ 7 Future Work
Future research could focus on refining the search criteria for the optimal model head by selecting an appropriate $\beta$ threshold. Additionally, a robust theoretical underpinning is needed to explain why shared representations improve the classification of known classes under distribution shifts. Integrating test-time or model-free domain adaptation methods, such as those by Wang et al. (2020), Zhang et al. (2023), or Saito et al. (2018), could further enhance OSDA performance. It would be valuable to compare our approach with large foundation models. Although curating such a benchmark would require access to publicly available pretraining datasets like LAION, which would require significant effort and present unforeseen challenges.
@@ Acknowledgments
We acknowledge the guidance of Dr. Rama Chellappa in this project. Furthermore, we acknowledge the support from the DARPA TIAMAT program under Grant No. HR00112490422. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the U.S. Government.
@@ A.1 Theoretical analysis
For a Gaussian two-feature PU problem, let $\hat{\bm{\eta}}=\bm{\eta}/\|\bm{\eta}\|$ and $\hat{\bm{\mu}}$ accordingly. We have that the AU-ROC of a model ${\mathbf{w}}\in{\mathbb{R}^{d}}$ is $Q(\frac{\langle{\mathbf{w}},\bm{\eta}-\bm{\mu}\rangle}{\sigma\sqrt{(2\|{\mathbf{w}}\|^{2}-\langle{\mathbf{w}},\hat{\bm{\eta}}\rangle^{2}-\langle{\mathbf{w}},\hat{\bm{\mu}}\rangle^{2})}})$.
@@
$=P(\langle{\mathbf{w}},-\bm{\eta}+\xi_{\mathrm{novel}}\rangle<\langle{\mathbf{w}},-\bm{\mu}+\xi_{\mathrm{{\mathcal{T}},0}}\rangle)$
$=P(\langle{\mathbf{w}},-\bm{\eta}+\bm{\mu}\rangle<\langle{\mathbf{w}},-\xi_{\mathrm{novel}}+\xi_{\mathrm{{\mathcal{T}},0}}\rangle),$
where $\xi_{\mathrm{novel}}\sim\mathcal{N}(0,\sigma^{2}(I_{d}-\hat{\bm{\mu}}\hat{\bm{\mu}}^{\top}))$ and $\xi_{{\mathcal{T}},0}\sim\mathcal{N}(0,\sigma^{2}(I_{d}-\hat{\bm{\eta}}\hat{\bm{\eta}}^{\top}))$. Then we have that $\langle{\mathbf{w}},\xi_{\mathrm{novel}}-\xi_{{\mathcal{T}},0}\rangle\sim\mathcal{N}\left(0,\sigma^{2}\left(2\|{\mathbf{w}}\|^{2}-\langle{\mathbf{w}},\hat{\bm{\eta}}\rangle^{2}-\langle{\mathbf{w}},\hat{\bm{\mu}}\rangle^{2}\right)\right)$. Dividing both sides of the argument above by a factor $\sigma\sqrt{2\|w\|^{2}-\langle{\mathbf{w}},\hat{\bm{\eta}}\rangle^{2}-\langle{\mathbf{w}},\hat{\bm{\mu}}\rangle^{2}}$ we get that for a variable $\xi\sim\mathcal{N}(0,1)$
$\mathrm{AU-ROC}({\mathbf{w}})$ $=P\left(\frac{\langle{\mathbf{w}},\bm{\mu}-\bm{\eta}\rangle}{\sigma\sqrt{2\|{\mathbf{w}}\|^{2}-\langle{\mathbf{w}},\hat{\bm{\eta}}\rangle^{2}-\langle w,\hat{\bm{\mu}}\rangle^{2}}}<\xi\right)$
$=Q\left(\frac{\langle{\mathbf{w}},\bm{\mu}-\bm{\eta}\rangle}{\sigma\sqrt{2\|{\mathbf{w}}\|^{2}-\langle{\mathbf{w}},\hat{\bm{\eta}}\rangle^{2}-\langle{\mathbf{w}},\hat{\bm{\mu}}\rangle^{2}}}\right)$
Note that when $\langle{\mathbf{w}},\bm{\mu}-\bm{\eta}\rangle>0$ then the AU-ROC is lower than $0.5$, and when $\langle{\mathbf{w}},\bm{\mu}-\bm{\eta}\rangle<0$ then $Q\left(\frac{\langle{\mathbf{w}},\bm{\mu}-\bm{\eta}\rangle}{\sigma\sqrt{2\|{\mathbf{w}}\|^{2}-\langle{\mathbf{w}},\hat{\bm{\eta}}\rangle^{2}-\langle{\mathbf{w}},\hat{\bm{\mu}}\rangle^{2}}}\right)\geq Q\left(\frac{\langle{\mathbf{w}},\bm{\mu}-\bm{\eta}\rangle}{\sqrt{2}\sigma\|{\mathbf{w}}\|}\right)$. Hence the RHS is a lower bound on the AU-ROC.
@@
$C_{d}\log(\frac{c_{d}}{\delta})$ $\leq\frac{\sqrt{d}}{4\sqrt{N_{{\mathcal{S}}}+N_{{\mathcal{T}}}}},$
$r_{\mu}$ $\geq\frac{1}{2\sqrt{N_{{\mathcal{S}}}+N_{{\mathcal{T}}}}},$
$r_{\eta}$ $\leq\frac{r_{\mu}}{4+8r_{\mu}\sqrt{N_{{\mathcal{T}},1}}}-\frac{1}{\sqrt{N_{{\mathcal{S}}}+N_{{\mathcal{T}}}}}$
There exists $C_{d},c_{d}$ such that for any $\delta$, if we fix $\bm{\mu},\bm{\eta},r_{\mu},r_{\eta},N_{{\mathcal{S}}},N_{{\mathcal{T}},0},N_{{\mathcal{T}},1}$ that are in the above set, and drawing training data $(S_{\mathcal{S}},S_{\mathcal{T}})$ as described in problem definition. Let ${\mathbf{w}}_{\mathrm{\text{DD}}}$ as defined in eq.˜2, then with probability at least $1-\delta$ we have that $AU-ROC({\mathbf{w}}_{\mathrm{\text{DD}}})\leq 0.5$.
@@
$\mathcal{L}({\mathbf{w}},\lambda,\nu)=\|{\mathbf{w}}\|^{2}+\lambda_{{\mathcal{S}}}^{\top}(\mathbf{1}_{{\mathcal{S}}}-\bm{X}_{{\mathcal{S}}}{\mathbf{w}})+\lambda_{{\mathcal{T}}}^{\top}(\mathbf{1}_{{\mathcal{T}}}+\bm{X}_{{\mathcal{T}}}{\mathbf{w}})+\nu\cdot({\mathbf{w}}^{\top}(\bm{\mu}-\bm{\eta}))$
@@
$\nabla_{{\mathbf{w}}}\mathcal{L}({\mathbf{w}},\lambda,\nu)=2{\mathbf{w}}-\bm{X}^{\top}_{{\mathcal{S}}}\lambda_{{\mathcal{S}}}+\bm{X}^{\top}_{{\mathcal{T}}}\lambda_{{\mathcal{T}}}+\nu\cdot(\bm{\mu}-\bm{\eta})=0$
$\Leftrightarrow{\mathbf{w}}=\frac{1}{2}\left(\nu\cdot(\bm{\eta}-\bm{\mu})+\bm{X}^{\top}_{{\mathcal{S}}}\lambda_{{\mathcal{S}}}-\bm{X}^{\top}_{{\mathcal{T}}}\lambda_{{\mathcal{T}}}\right)$
@@
$\mathcal{L}(\lambda,\nu)=\lambda_{{\mathcal{S}}}^{\top}\mathbf{1}_{{\mathcal{S}}}+\lambda_{{\mathcal{T}}}^{\top}\mathbf{1}_{{\mathcal{T}}}-\frac{1}{4}\|\nu\cdot(\bm{\eta}-\bm{\mu})+\bm{Z}^{\top}\lambda\|^{2}$
@@
Let $\alpha>0$, to be determined later, $\nu=\frac{(|U|-k)2\alpha}{|U|}$, and define the set $U=\mathcal{I}_{{\mathcal{T}},1}\cup\mathcal{I}_{\mathrm{Rand}}$ where $\mathcal{I}_{\mathrm{Rand}}$ has $\frac{|U|}{2\alpha}\nu$ indices randomly drawn from $\mathcal{I}_{{\mathcal{S}}}$ and half from $\mathcal{I}_{{\mathcal{T}},0}$. Hence, $|U|=N$ Set $\lambda=\frac{\alpha}{|U|}\mathbf{1}_{U}$ and plug back into the Lagrangian. We have that
$\bm{Z}^{\top}\lambda=\nu\bm{\mu}+\frac{\alpha}{|U|}\left[k\bm{\eta}+\sum_{i\in{U}}{\xi_{i}y_{i}}\right].$
@@
$\mathcal{L}(\alpha,\nu)=\alpha-\frac{1}{4}\|\bm{\eta}(\nu+k\frac{\alpha}{|U|})+\alpha\bar{\xi}_{U}\|^{2}$
@@
$\mathcal{L}(\alpha)=\alpha-\frac{1}{4}\|\bm{\eta}(\frac{\alpha(2|U|-k)}{|U|})+\alpha\bar{\xi}_{U}\|^{2}$
@@
$\|{\mathbf{w}}^{*}\|^{2}\geq\mathcal{L}(\alpha^{*})$ $=4\|\bm{\eta}(\frac{2|U|-k}{|U|})+\bar{\xi}_{U}\|^{-2}$
$=4\left(r^{2}_{\eta}(2-k/|U|)^{2}+\|\bar{\xi}_{U}\|^{2}+2(2-k/|U|)\bm{\eta}^{\top}\bar{\xi}_{U}\right)^{-1}$
@@
${\mathbf{w}}_{\mathrm{\text{DD}}}=2\left(r^{2}_{\mu}-\sigma r_{\mu}t\right)^{-1}\cdot\bm{\mu}-4\sigma^{-2}kd^{-1}\cdot\bar{\xi}_{{\mathcal{T}},1},$
satisfies $y_{i}{\mathbf{w}}_{\mathrm{\text{DD}}}^{\top}{\mathbf{x}}_{i}\geq 1$ for all $i\in{U}$ and ${\mathbf{w}}_{\mathrm{\text{DD}}}^{\top}(\bm{\mu}-\bm{\eta})\geq 0$.